# The Exact Mixing Time for Trees of Fixed Order and Diameter

### v1

@@ 5 Conclusion
There are many open questions about random walks on the family $\mathcal{T}_{n,d}$ of trees of order $n$ and diameter $d$. Two of the current authors have shown that the balanced double broom also achieves the maximum value of the reset time $\mathsf{T}_{\mathsf{reset}}=\sum_{v\in V}\pi_{v}H(v,\pi)$, see [2]. Another quantity of interest is the best mixing time $\mathsf{T}_{\mathsf{bestmix}}=\min_{v\in V}H(v,\pi)$.
It would also be interesting to identify the minimizing structure in $\mathcal{T}_{n,d}$ for each of these exact mixing measures. Finally, we could ask these same questions about the family of all graphs of diameter $d$: which graphs achieve the extremal values of each of these mixing measures?

### v2

@@ 5 Conclusion
There are many open questions about random walks on the family $\mathcal{T}_{n,d}$ of trees of order $n$ and diameter $d$. For example, it would be interesting to characterize the behavior of the reset time $\mathsf{T}_{\mathsf{reset}}=\sum_{v\in V}\pi_{v}H(v,\pi)$ and thee best mixing time $\mathsf{T}_{\mathsf{bestmix}}=\min_{v\in V}H(v,\pi)$. Another direction would be to identify the minimizing structure in $\mathcal{T}_{n,d}$ for each of these exact mixing measures. Finally, we could ask these same questions about the family of all graphs of diameter $d$: which graphs achieve the extremal values of each of these mixing measures?